# Proper Learnability and the Role of Unlabeled Data

### v1

@@ Abstract
We complement this with impossibility results which obstruct any characterization of proper learnability in the classic (realizable) PAC model. First, we observe that there are problems whose proper learnability is logically undecidable, i.e., independent of the ZFC axioms. We then show that proper learnability is not a monotone property of the underlying hypothesis class, and that it is not a local property (in a precise sense). We also point out how the non-monotonicity of proper learning obstructs relaxations of the distribution-fixed model that preserve proper learnability, including natural notions of class-conditional learning of the unlabeled data distribution. Our impossibility results all hold even for the fundamental setting of multiclass classification, and go through a reduction of EMX learning (Ben-David et al., 2019) to proper classification which may be of independent interest.
@@ 1 Introduction
Multiclass classification proceeds identically as in binary classification, save for the fact that the collection of possible labels for the data — denoted $\mathcal{Y}$ — is permitted to be arbitrarily large, perhaps even infinite. To what extent do insights from binary classification extend to the multiclass case? Perhaps less than one may expect. Daniely and Shalev-Shwartz (2014) showed that there exist multiclass classification problems which are learnable yet cannot be learned by any proper learner. Notably, this demonstrates that ERM, perhaps the quintessential workhorse of machine learning (and particularly binary classification), does not enjoy the same success in multiclass classification.
In fact, the task of characterizing multiclass learnability via some choice of dimension, analogous to the VC dimension for binary classification, remained a major open problem for decades. Brukhim et al. (2022) recently demonstrated in a breakthrough result that learnability is in fact characterized by the Daniely-Shalev-Shwartz (DS) dimension. Further, they exhibited learners for all classes of finite DS dimension, based upon certain extensions of the one-inclusion graph predictor of Haussler et al. (1994) and several novel ideas such as list PAC learning. Interestingly, the learner of Brukhim et al. (2022) employs techniques which are strikingly different from — and more intricate than — ERM and the standard algorithmic approaches of binary classification. The complexity of existing (improper) learners for multiclass classification, and the necessity of some such complexity by the result of Daniely and Shalev-Shwartz (2014), raises a natural question: Can one succeed with simpler learning rules, under additional assumptions on $\mathcal{H}$ or on the learning model?
@@
In light of our positive result, it is natural to ask whether one can draw a connection between proper learnability and unsupervised learning (i.e., the ability to infer distributional information about the unlabeled data distribution $\mathcal{D}$) in the classic PAC model. Perhaps $\mathcal{H}$ is properly learnable precisely when $\mathcal{D}$ can be learned in some “class-conditional” sense which depends upon $\mathcal{H}$? Several such conditions have been proposed by Hopkins et al. (2023) to study binary classification, including Weak TV-learning, Strong TV-learning, and Exact TV-learning. All such definitions are monotone, however, and thus — by our previous impossibility result — cannot characterize proper learnability. More generally, any notion of learning the marginal in a class-condition manner will likely take the form of a monotone property, and thus fail to characterize proper learnability. In short, a precise characterization of proper learnability may require a fundamentally different approach than the standard techniques of (improper) supervised learning.
@@ 1.1 Related Work
We focus primarily — but not exclusively — on the setting of multiclass classification, i.e., learning under the 0-1 loss function. When $|\mathcal{Y}|=2$, one recovers binary classification, for which learnability is characterized by the VC dimension and empirical risk minimization (ERM) is a nearly-optimal learner (Blumer et al., 1989; Shalev-Shwartz and Ben-David, 2014). As ERM is proper, learnability is thus equivalent to proper learnability in the binary case. [^1] Attaining the optimal sample complexity, however, is known to require improper learning in general. Interestingly, recent work has demonstrated that the improperness requirement for optimal learning can be satisfied using simple aggregations of proper learners, such as a majority of only 3 ERM learners (Hanneke, 2016; Larsen, 2023; Aden-Ali et al., 2024). In the multiclass setting, however, there are learnable classes which cannot be learned by any aggregation of a finite number of proper learners (Asilis et al., 2025). In the multiclass case, $\mathcal{Y}$ is permitted to be of arbitrarily large size (even infinite), and the equivalence between proper learnability and improper learnability from the binary case was shown to fail by Daniely and Shalev-Shwartz (2014). They also proposed the Daniely-Shalev-Shwartz (DS) dimension, and conjectured that it characterizes improper multiclass learnability. This was recently confirmed in a breakthrough result of Brukhim et al. (2022), resolving a long-standing open question. Regarding algorithmic templates for multiclass learning, relatively little is known: Brukhim et al. (2022) designed one learner for general DS classes, using an intricate sequence of arguments and algorithmic techniques (e.g., list PAC learning, sample compression, one-inclusion graphs, etc.). It is natural to ask for simpler learners than that of Brukhim et al. (2022), perhaps which bear a closer resemblance to algorithms enjoying practical success (e.g., structural risk minimization (SRM)). Recently, Asilis et al. (2024b) made some progress by demonstrating that there always exist optimal learners taking the form of unsupervised local regularization, a certain relaxation of classical regularization. The proof is non-constructive, however, saying little about the precise form of the regularizer or the learner. Perhaps most relevant to our work is the line of research studying learnability via ERM. This includes work demonstrating that there can be arbitrarily large gaps between the sample complexities of different ERM learners, and that the sample complexity of ERM is closely related to the graph dimension (Daniely et al., 2015). Learnability by any ERM learner is not equivalent to proper learnability, however, and thus this does not directly address our primary question. Regarding the issue of optimal proper learning, Bousquet et al. (2020) studied the conditions under which a binary hypothesis class $\mathcal{H}$ can be learned with optimal sample complexity by a proper learner, and established a characterization via finiteness of the dual Helly number, under general conditions on $\mathcal{H}$.
@@
There is a long line of work studying the power of unlabeled data in learning, often formalized by the setting in which a learner receives both labeled and unlabeled datapoints, i.e., semi-supervised learning (SSL) (Kääriäinen, 2005; Zhu, 2005; Chapelle et al., 2006; Van Engelen and Hoos, 2020). One direction has studied SSL under the assumption that there is a relationship between the unlabeled data distribution $\mathcal{D}$ and the true labeling function $h^{*}$, and demonstrated results supporting the power of unlabeled data in this setting (Castelli and Cover, 1995; Seeger, 2000; Rigollet, 2007; Singh et al., 2008; Niyogi, 2013). Another line of work demonstrates that in that absence of any such assumptions, unlabeled data has little effect in binary classification from a worst-case perspective (Ben-David et al., 2008; Darnstädt and Simon, 2011; Göpfert et al., 2019). Yet another direction of work studies the power of unlabeled training points from a fine-grained perspective, examining learners’ sample complexities on particular data distributions rather than on a worst-case basis, and establishes the value of unlabeled data in learning binary classes of infinite VC dimension (Darnstädt et al., 2013). We study a setting which makes no assumption on the unlabeled distribution $\mathcal{D}$ or the true labeling function $h^{*}\in\mathcal{H}$, but assumes that the learner receives complete information of $\mathcal{D}$. The learner is then judged on a worst-case basis over all possible (realizable) distributions. This most closely aligns with the “utopian” model of SSL studied by Ben-David et al. (2008) and Lu (2009). Notably, Göpfert et al. (2019) demonstrated that this setting — which they refer to as simply “knowing the marginal” — is of no additional help for binary classification. (I.e., the worst-case expected error rate of a learner does not improve by granting it knowledge of the marginal.) Our analogous result can be seen as extending this finding to a broader collection of bounded metric loss functions (Theorem 3.4).
@@
In Section 4, we establish several obstructions to characterizing proper learnability in multiclass classification, including by demonstrating that there exist classes $\mathcal{H}$ for which it is logically undecidable whether $\mathcal{H}$ can be properly learned. That is, within the ZFC axioms it can be neither proven nor disproven that $\mathcal{H}$ is properly learnable. This result builds upon the breakthrough work of Ben-David et al. (2019), which established that the learnability of certain EMX (Estimating the Maximum) learning problems can be undecidable. Notably, Hanneke and Yang (2023) established an equivalence between certain EMX learning problems and bandit problems in order to establish that bandit learnability can likewise be undecidable. A related line of work, also inspired by Ben-David et al. (2019), investigates the algorithmic decidability of learning, i.e., examining whether problems can be learned using learners which are computable, rather than merely abstract mathematical functions (Agarwal et al., 2020). Recent developments in this area have established that there exist VC classes which cannot be learned by any computable learner (Sterkenburg, 2022), and that learnability via (improper) computable learners is instead characterized by the effective VC dimension, which roughly measures the smallest cardinality $k$ for which one can always compute a behavior on any $k+1$ distinct unlabeled points which $\mathcal{H}$ cannot express (Delle Rose et al., 2023). Notably, this characterization holds for binary classification over the domain of the natural numbers; for binary classification over more general computable metric spaces, see Ackerman et al. (2022). Further work includes that of Hasrati and Ben-David (2023) on computable online learning, Gourdeau et al. (2024) on the computability of robust PAC learning, and Caro (2023), which studies the computability of learning when learners are equipped with a restricted form of black-box access to the underlying hypothesis class $\mathcal{H}$.
@@ 2.1 Notation
For a natural number $n\in\mathbb{N}$, $[n]$ denotes the set $\{1,2,\ldots,n\}$. When $M$ is a measurable space, we let $\Delta(M)$ denote the collection of all probability measures over $M$. Finite sets $M_{\mathrm{fin}}$ are thought of as measurable spaces by endowing them with the discrete $\sigma$-algebra by default. $\operatorname{Unif}(M_{\mathrm{fin}})$ denotes the uniform distribution over $M_{\mathrm{fin}}$.
@@ 2.2 Learning Theory
A successful learner is one which attains vanishingly small error when trained on increasingly large datasets, as formalized by Valiant’s celebrated Probably Approximately Correct (PAC) learning model (Valiant, 1984).
@@
In the realizable setting, SRM learners are sometimes defined as those which minimize the regularization value $\psi(h)$ subject to a hard constraint on attaining zero training error (Asilis et al., 2024a). This perspective is essentially equivalent to Definition 2.5, as one can normalize $\psi$ to have output strictly less than $\frac{1}{|S|}$ for the case of classification. (Note that this normalization depends upon $|S|$, however.)
@@ 3 Proper Learning Through Distributional Regularization
The distribution-fixed model reflects the setting in which the marginal distribution over unlabeled data, $\mathcal{D}$, is fully known to the learner at training time, yet the learner is judged on its worst-case performance across any choice of $\mathcal{D}$ (and true labeling function $h^{*}\in\mathcal{H}$). Other models, including the seminal work of Benedek and Itai (1991), focus on the case in which the marginal $\mathcal{D}$ does not vary (and only the true labeling function can vary). In this setting, the learner is endowed with complete information of $\mathcal{D}$ “by default”, i.e., because its performance is only examined on distributions which share this marginal. In short, Benedek and Itai (1991) adopt an instance-optimal perspective on learning under a particular marginal distribution. The version we study can be thought of as intermediate between the classical PAC model and that of Benedek and Itai. Informally, we maintain a worst-case perspective but equip learners with a complete understanding of the unlabeled data.
@@
The first inequality makes use of a standard repetition argument; a learner incurring expected error at most $\epsilon/2$ can be repeatedly trained on separate datasets and tested on a validation set in order to attain a high-probability guarantee. The second inequality invokes the claim whose proof we deferred to Section A.1. The third inequality follows immediately from the fact that the loss function is bounded above by 1. Conclude by noting that $114e\approx 10.87<$ 11. ∎
@@
Let $\mathcal{X}$ be a finite domain, $\mathcal{Y}$ a finite label space, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Let $Q\in\Delta(\mathcal{H})$ be a distribution over $\mathcal{H}$, and $A$ be a Bayesian learner with respect to $Q$. That is, upon receiving a training sample $S$, $A$ emits $Q|S\in\Delta(\mathcal{H})$, the restriction of $D$ to those $h\in\mathcal{H}$ with $L_{S}(h)=0$. Then $A$ is a distributional SRM learner.
@@
Then an SRM learner induced by $\psi$ is tasked with outputing a distribution $P$ which minimizes empirical error (i.e., is supported on $L_{S}^{-1}(0)$) while minimizing relative entropy to $Q$. By a standard result, the distribution supported on $L_{S}^{-1}(0)$ with minimal relative entropy to $Q$ is precisely $Q|S$, the restriction of $Q$ to $L_{S}^{-1}(0)$. (See, e.g., Asilis et al. (2024b, Lemma 55).) This completes the argument. ∎
@@
We now argue that the expected error incurred by $A_{\text{prop}}$ is optimal, up to a factor of 2. To this end, fix a test point $x\in\mathcal{X}$ and let $P=\{p_{y}\}_{y\in\mathcal{Y}}$ denote the posterior distribution over the true label at $x$ (upon conditioning $\Lambda$ by $S$). By linearity of expectation, the optimal prediction at $x$ can be assumed to be deterministic, i.e., predicting a fixed label $\hat{y}\in\mathcal{Y}$. Thus, the expected error incurred by the optimal learner at test point $x$ is $\operatorname*{\mathbb{E}}_{y\sim P}\ell(y,\hat{y})$. The error incurred by $A_{\text{prop}}$, in contrast, is $\operatorname*{\mathbb{E}}_{y,y^{\prime}\sim P}\ell(y,y^{\prime})$. With an application of the triangle inequality, we have:
@@
Our proof of Theorem 3.9 bears similarities to the proof of Darnstädt and Simon (2011, Lemma 5), which also models learning with respect to a fixed marginal distribution as a zero-sum game. Though the goals of their paper and this particular lemma are substantially different from ours, their analysis of the learner’s best response problem is conceptually similar. One key difference is that their proof exploits structure particular to binary classification, which is their focus. Another difference is that they restrict attention to proper learning out of the gate, as this is without loss for binary classification, whereas we allow improper learning and conclude that properness is without much loss for more general problem classes. Finally, we aggregate the (Bayesian, and proper) best responses of the learner into a near-optimal learner of the same desired form, departing from the concerns and technical approach of their paper.
@@
It is natural to ask whether the conclusion of Theorem 3.9 can be achieved by assuming a “softer” form of distribution-fixed learning. For instance, perhaps one can remain in the classic PAC model yet assume that $\mathcal{H}$ is sufficiently simple such that a learner $A$ can use the unlabeled data in $S$ in order to learn the marginal distribution $\mathcal{D}$ over $\mathcal{X}$ in some “class-conditional” sense. (E.g., to learn $\mathcal{D}$ sufficiently well so as to estimate $L_{\mathcal{D}_{h}}(h^{\prime})$ for all pairs $(h,h^{\prime})\in\mathcal{H}^{2}$.) This line of inquiry is studied by Hopkins et al. (2023) for binary classification in the distribution-family model, in which the marginal distribution $\mathcal{D}$ over $\mathcal{X}$ is restricted to a certain collection of distributions at the outset. They introduce precisely such “class-conditional” notions of learning the marginal $\mathcal{D}$, and provide distinct necessary and sufficient conditions for PAC learnability based upon learnability of $\mathcal{D}$.
Notably, however, the difficulty in Hopkins et al. (2023) arises from studying the distribution-family model, and the complexity which it can endow binary classification. Further, Hopkins et al. (2023) do not emphasize the particular algorithmic or structural form of the learner, as we do. In PAC learning with respect to all realizable distributions — as we study — binary classification problems are also well-known to be learnable by proper learners whenever learning is possible.
It may be natural to ask, then, whether the techniques of Hopkins et al. (2023) are applicable for classical PAC learning beyond the binary setting. To this, we present a negative result in Section 4, by demonstrating that proper learnability is not a monotone property. That is, there exist hypothesis classes $\mathcal{H}_{0}\subsetneq\mathcal{H}_{1}\subsetneq\mathcal{H}_{2}$ in multiclass classification such that only $\mathcal{H}_{0}$ and $\mathcal{H}_{2}$ are properly learnable. All notions of class-conditional learnability introduced by Hopkins et al. (2023), however, are monotone (e.g., Weak TV-learning, Strong TV-learning, Exact TV-learning). As such, any natural weakening of the distribution-fixed assumption — reflecting the ability to learn $\mathcal{D}$ in a “class-conditional” way — is unlikely to characterize proper learnability.
@@
A central point of Theorem 3.9 is that the proper learner is optimal in terms of its expected error, up to a constant factor of 2. In particular, even for finite problems in classical PAC learning (i.e., finite $\mathcal{X}$ and $\mathcal{Y}$), there are known to be problems exhibiting arbitrarily large gaps in sample complexity between proper and improper learners. (See Daniely and Shalev-Shwartz (2014, Theorem 1), along with the compactness result of Asilis et al. (2024c) which equates the sample complexity of a learner to its worst-case over finite subproblems.) As such, Theorem 3.9 would not hold when stated in the classic PAC model. Furthermore, Theorem 3.9 establishes an equivalence between proper learnability and learnability by (distributional) SRM in the distribution-fixed model. It is natural to ask whether such an equivalence might hold more generally, such as in the classic PAC model. We conjecture that it may indeed be so, at least for the case of multiclass classification.

### v2

@@ Abstract
We complement this with impossibility results which obstruct any characterization of proper learnability in the classic (realizable) PAC model. First, we observe that there are problems whose proper learnability is logically undecidable, i.e., independent of the ZFC axioms. We then show that proper learnability is not a monotone property of the underlying hypothesis class, and that it is not a local property (in a precise sense). We also point out how the non-monotonicity of proper learning obstructs relaxations of the distribution-fixed model that preserve proper learnability, including natural notions of class-conditional learning of the unlabeled data distribution. Our impossibility results all hold even for the fundamental setting of multiclass classification, and go through a reduction of EMX learning (ben2019learnability) to proper classification which may be of independent interest.
@@ 1 Introduction
Multiclass classification proceeds identically as in binary classification, save for the fact that the collection of possible labels for the data — denoted $\mathcal{Y}$ — is permitted to be arbitrarily large, perhaps even infinite. To what extent do insights from binary classification extend to the multiclass case? Perhaps less than one may expect. DS14 showed that there exist multiclass classification problems which are learnable yet cannot be learned by any proper learner. Notably, this demonstrates that ERM, perhaps the quintessential workhorse of machine learning (and particularly binary classification), does not enjoy the same success in multiclass classification.
In fact, the task of characterizing multiclass learnability via some choice of dimension, analogous to the VC dimension for binary classification, remained a major open problem for decades. brukhim2022characterization recently demonstrated in a breakthrough result that learnability is in fact characterized by the Daniely-Shalev-Shwartz (DS) dimension. Further, they exhibited learners for all classes of finite DS dimension, based upon certain extensions of the one-inclusion graph predictor of haussler1994predicting and several novel ideas such as list PAC learning. Interestingly, the learner of brukhim2022characterization employs techniques which are strikingly different from — and more intricate than — ERM and the standard algorithmic approaches of binary classification. The complexity of existing (improper) learners for multiclass classification, and the necessity of some such complexity by the result of DS14, raises a natural question: Can one succeed with simpler learning rules, under additional assumptions on $\mathcal{H}$ or on the learning model?
@@
In light of our positive result, it is natural to ask whether one can draw a connection between proper learnability and unsupervised learning (i.e., the ability to infer distributional information about the unlabeled data distribution $\mathcal{D}$) in the classic PAC model. Perhaps $\mathcal{H}$ is properly learnable precisely when $\mathcal{D}$ can be learned in some “class-conditional” sense which depends upon $\mathcal{H}$? Several such conditions have been proposed by hopkins2023pac to study binary classification, including Weak TV-learning, Strong TV-learning, and Exact TV-learning. All such definitions are monotone, however, and thus — by our previous impossibility result — cannot characterize proper learnability. More generally, any notion of learning the marginal in a class-condition manner will likely take the form of a monotone property, and thus fail to characterize proper learnability. In short, a precise characterization of proper learnability may require a fundamentally different approach than the standard techniques of (improper) supervised learning.
@@ 1.1 Related Work
We focus primarily — but not exclusively — on the setting of multiclass classification, i.e., learning under the 0-1 loss function. When $|\mathcal{Y}|=2$, one recovers binary classification, for which learnability is characterized by the VC dimension and empirical risk minimization (ERM) is a nearly-optimal learner (BEHW89; shalev2014understanding). As ERM is proper, learnability is thus equivalent to proper learnability in the binary case. [^1] Attaining the optimal sample complexity, however, is known to require improper learning in general. Interestingly, recent work has demonstrated that the improperness requirement for optimal learning can be satisfied using simple aggregations of proper learners, such as a majority of only 3 ERM learners (hanneke2016optimal; larsen2023bagging; aden2024majority). In the multiclass setting, however, there are learnable classes which cannot be learned by any aggregation of a finite number of proper learners (asilisunderstanding). In the multiclass case, $\mathcal{Y}$ is permitted to be of arbitrarily large size (even infinite), and the equivalence between proper learnability and improper learnability from the binary case was shown to fail by DS14. They also proposed the Daniely-Shalev-Shwartz (DS) dimension, and conjectured that it characterizes improper multiclass learnability. This was recently confirmed in a breakthrough result of brukhim2022characterization, resolving a long-standing open question. Regarding algorithmic templates for multiclass learning, relatively little is known: brukhim2022characterization designed one learner for general DS classes, using an intricate sequence of arguments and algorithmic techniques (e.g., list PAC learning, sample compression, one-inclusion graphs, etc.). It is natural to ask for simpler learners than that of brukhim2022characterization, perhaps which bear a closer resemblance to algorithms enjoying practical success (e.g., structural risk minimization (SRM)). Recently, asilis2024regularization made some progress by demonstrating that there always exist optimal learners taking the form of unsupervised local regularization, a certain relaxation of classical regularization. The proof is non-constructive, however, saying little about the precise form of the regularizer or the learner. Perhaps most relevant to our work is the line of research studying learnability via ERM. This includes work demonstrating that there can be arbitrarily large gaps between the sample complexities of different ERM learners, and that the sample complexity of ERM is closely related to the graph dimension (daniely2015multiclass). Learnability by any ERM learner is not equivalent to proper learnability, however, and thus this does not directly address our primary question. Regarding the issue of optimal proper learning, bousquet2020proper studied the conditions under which a binary hypothesis class $\mathcal{H}$ can be learned with optimal sample complexity by a proper learner, and established a characterization via finiteness of the dual Helly number, under general conditions on $\mathcal{H}$.
@@
There is a long line of work studying the power of unlabeled data in learning, often formalized by the setting in which a learner receives both labeled and unlabeled datapoints, i.e., semi-supervised learning (SSL) (kaariainen2005generalization; zhu2005semi; chapelle2006semi; van2020survey). One direction has studied SSL under the assumption that there is a relationship between the unlabeled data distribution $\mathcal{D}$ and the true labeling function $h^{*}$, and demonstrated results supporting the power of unlabeled data in this setting (castelli1995exponential; seeger2000learning; balcan2005pac; rigollet2007generalization; singh2008unlabeled; balcan2010semi; niyogi2013manifold). Another line of work demonstrates that in that absence of any such assumptions, unlabeled data has little effect in binary classification from a worst-case perspective (ben2008does; darnstadt2011smart; gopfert2019can). Yet another direction of work studies the power of unlabeled training points from a fine-grained perspective, examining learners’ sample complexities on particular data distributions rather than on a worst-case basis, and establishes the value of unlabeled data in learning binary classes of infinite VC dimension (darnstadt2013unlabeled). We study a setting which makes no assumption on the unlabeled distribution $\mathcal{D}$ or the true labeling function $h^{*}\in\mathcal{H}$, but assumes that the learner receives complete information of $\mathcal{D}$. The learner is then judged on a worst-case basis over all possible (realizable) distributions. This most closely aligns with the “utopian” model of SSL studied by ben2008does and lu2009fundamental. Notably, gopfert2019can demonstrated that this setting — which they refer to as simply “knowing the marginal” — is of no additional help for binary classification. (I.e., the worst-case expected error rate of a learner does not improve by granting it knowledge of the marginal.) Our analogous result can be seen as extending this finding to a broader collection of bounded metric loss functions (Theorem 3.4).
@@
In Section 4, we establish several obstructions to characterizing proper learnability in multiclass classification, including by demonstrating that there exist classes $\mathcal{H}$ for which it is logically undecidable whether $\mathcal{H}$ can be properly learned. That is, within the ZFC axioms it can be neither proven nor disproven that $\mathcal{H}$ is properly learnable. This result builds upon the breakthrough work of ben2019learnability, which established that the learnability of certain EMX (Estimating the Maximum) learning problems can be undecidable. Notably, bandit-undecidable established an equivalence between certain EMX learning problems and bandit problems in order to establish that bandit learnability can likewise be undecidable. A related line of work, also inspired by ben2019learnability, investigates the algorithmic decidability of learning, i.e., examining whether problems can be learned using learners which are computable, rather than merely abstract mathematical functions (agarwal2020learnability). Recent developments in this area have established that there exist VC classes which cannot be learned by any computable learner (sterkenburg2022characterizations), and that learnability via (improper) computable learners is instead characterized by the effective VC dimension, which roughly measures the smallest cardinality $k$ for which one can always compute a behavior on any $k+1$ distinct unlabeled points which $\mathcal{H}$ cannot express (delle2023find). Notably, this characterization holds for binary classification over the domain of the natural numbers; for binary classification over more general computable metric spaces, see ackerman2022computable. Further work includes that of hasrati2023computable on computable online learning, gourdeau2024computability on the computability of robust PAC learning, and caro2023undecidability, which studies the computability of learning when learners are equipped with a restricted form of black-box access to the underlying hypothesis class $\mathcal{H}$.
@@ 2.1 Notation
For a natural number $\in\mathbb{N}, [n][n$] denotes the set $\{1,2,\ldots,n\}$. When $M$ is a measurable space, we let $\Delta(M)$ denote the collection of all probability measures over $M$. Finite sets $M_{\mathrm{fin}}$ are thought of as measurable spaces by endowing them with the discrete $\sigma$-algebra by default. $\operatorname{Unif}(M_{\mathrm{fin}})$ denotes the uniform distribution over $M_{\mathrm{fin}}$.
@@ 2.2 Learning Theory
A successful learner is one which attains vanishingly small error when trained on increasingly large datasets, as formalized by Valiant’s celebrated Probably Approximately Correct (PAC) learning model (valiant1984theory).
@@
In the realizable setting, SRM learners are sometimes defined as those which minimize the regularization value $\psi(h)$ subject to a hard constraint on attaining zero training error (asilis2024open). This perspective is essentially equivalent to Definition 2.5, as one can normalize $\psi$ to have output strictly less than $\frac{1}{|S|}$ for the case of classification. (Note that this normalization depends upon $|S|$, however.)
@@ 3 Proper Learning Through Distributional Regularization
The distribution-fixed model reflects the setting in which the marginal distribution over unlabeled data, $\mathcal{D}$, is fully known to the learner at training time, yet the learner is judged on its worst-case performance across any choice of $\mathcal{D}$ (and true labeling function $h^{*}\in\mathcal{H}$). Other models, including the seminal work of benedek1991learnability, focus on the case in which the marginal $\mathcal{D}$ does not vary (and only the true labeling function can vary). In this setting, the learner is endowed with complete information of $\mathcal{D}$ “by default”, i.e., because its performance is only examined on distributions which share this marginal. In short, benedek1991learnability adopt an instance-optimal perspective on learning under a particular marginal distribution. The version we study can be thought of as intermediate between the classical PAC model and that of Benedek and Itai. Informally, we maintain a worst-case perspective but equip learners with a complete understanding of the unlabeled data.
@@
The first inequality makes use of a standard repetition argument; a learner incurring expected error at most $\epsilon/2$ can be repeatedly trained on separate datasets and tested on a validation set in order to attain a high-probability guarantee. The second inequality invokes the claim whose proof we deferred to Section A.1. The third inequality follows immediately from the fact that the loss function is bounded above by 1. Conclude by noting that $4e\approx 10.87<11$. ∎
@@
Let $\mathcal{X}$ be a finite domain, $\mathcal{Y}$ a finite label space, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Let $Q\in\Delta(\mathcal{H})$ be a distribution over $\mathcal{H}$, and $A$ be a Bayesian learner with respect to $Q$. That is, upon receiving a training sample $S$, AA emits $Q|S\in\Delta(\mathcal{H})$, the restriction of $D$ to those $h\in\mathcal{H}$ with $L_{S}(h)=0$. Then $A$ is a distributional SRM learner.
@@
Then an SRM learner induced by $\psi$ is tasked with outputing a distribution $P$ which minimizes empirical error (i.e., is supported on $L_{S}^{-1}(0)$) while minimizing relative entropy to $Q$. By a standard result, the distribution supported on $L_{S}^{-1}(0)$ with minimal relative entropy to $Q$ is precisely $Q|S$, the restriction of $Q$ to $L_{S}^{-1}(0)$. (See, e.g., asilis2024regularization.) This completes the argument. ∎
@@
[…]
@@
Our proof of Theorem 3.9 bears similarities to the proof of darnstadt2011smart, which also models learning with respect to a fixed marginal distribution as a zero-sum game. Though the goals of their paper and this particular lemma are substantially different from ours, their analysis of the learner’s best response problem is conceptually similar. One key difference is that their proof exploits structure particular to binary classification, which is their focus. Another difference is that they restrict attention to proper learning out of the gate, as this is without loss for binary classification, whereas we allow improper learning and conclude that properness is without much loss for more general problem classes. Finally, we aggregate the (Bayesian, and proper) best responses of the learner into a near-optimal learner of the same desired form, departing from the concerns and technical approach of their paper.
@@
It is natural to ask whether the conclusion of Theorem 3.9 can be achieved by assuming a “softer” form of distribution-fixed learning. For instance, perhaps one can remain in the classic PAC model yet assume that $\mathcal{H}$ is sufficiently simple such that a learner $A$ can use the unlabeled data in $S$ in order to learn the marginal distribution $\mathcal{D}$ over $\mathcal{X}$ in some “class-conditional” sense. (E.g., to learn $\mathcal{D}$ sufficiently well so as to estimate $L_{\mathcal{D}_{h}}(h^{\prime})$ for all pairs $(h,h^{\prime})\in\mathcal{H}^{2}$.) This line of inquiry is studied by hopkins2023pac for binary classification in the distribution-family model, in which the marginal distribution $\mathcal{D}$ over $\mathcal{X}$ is restricted to a certain collection of distributions at the outset. They introduce precisely such “class-conditional” notions of learning the marginal $\mathcal{D}$, and provide distinct necessary and sufficient conditions for PAC learnability based upon learnability of $\mathcal{D}$.
Notably, however, the difficulty in hopkins2023pac arises from studying the distribution-family model, and the complexity which it can endow binary classification. Further, hopkins2023pac do not emphasize the particular algorithmic or structural form of the learner, as we do. In PAC learning with respect to all realizable distributions — as we study — binary classification problems are also well-known to be learnable by proper learners whenever learning is possible.
It may be natural to ask, then, whether the techniques of hopkins2023pac are applicable for classical PAC learning beyond the binary setting. To this, we present a negative result in Section 4, by demonstrating that proper learnability is not a monotone property. That is, there exist hypothesis classes $\mathcal{H}_{0}\subsetneq\mathcal{H}_{1}\subsetneq\mathcal{H}_{2}$ in multiclass classification such that only $\mathcal{H}_{0}$ and $\mathcal{H}_{2}$ are properly learnable. All notions of class-conditional learnability introduced by hopkins2023pac, however, are monotone (e.g., Weak TV-learning, Strong TV-learning, Exact TV-learning). As such, any natural weakening of the distribution-fixed assumption — reflecting the ability to learn $\mathcal{D}$ in a “class-conditional” way — is unlikely to characterize proper learnability.
@@
A central point of Theorem 3.9 is that the proper learner is optimal in terms of its expected error, up to a constant factor of 2. In particular, even for finite problems in classical PAC learning (i.e., finite $\mathcal{X}$ and $\mathcal{Y}$), there are known to be problems exhibiting arbitrarily large gaps in sample complexity between proper and improper learners. (See DS14, along with the compactness result of asilis2024transductive which equates the sample complexity of a learner to its worst-case over finite subproblems.) As such, Theorem 3.9 would not hold when stated in the classic PAC model. Furthermore, Theorem 3.9 establishes an equivalence between proper learnability and learnability by (distributional) SRM in the distribution-fixed model. It is natural to ask whether such an equivalence might hold more generally, such as in the classic PAC model. We conjecture that it may indeed be so, at least for the case of multiclass classification.